OpenStudy (anonymous):

Which of the following is equivalent to the expression below? 1+sec(theta) ----------------- tan(theta)+sin(theta)

OpenStudy (anonymous):

\[\frac{1 + sec(\theta)}{tan(\theta) + sin(\theta)} = \frac{1 + \frac{1}{cos(\theta)}}{\frac{sin(\theta)}{cos(\theta)} + sin(\theta)}\]

OpenStudy (anonymous):

I dunno what options you have, these trig identities can get complicated.

OpenStudy (anonymous):

csc(theta) tan(theta) -1 sec(theta) cot(theta)

OpenStudy (anonymous):

Well, from the denominator you can factor out a sin(theta) from each term.. what do you have then?

OpenStudy (anonymous):

uh (1+1/cos(theta))/(cos(theta))?

OpenStudy (anonymous):

no. try factoring again \[\frac{sin\theta}{cos\theta} + sin\theta = ?\]

OpenStudy (anonymous):

sec(theta)

OpenStudy (anonymous):

that's the answer yes, but not what I was asking.. I assume you saw the solution after you factored it?

OpenStudy (anonymous):

err actually no, that's not the answer. nvm

OpenStudy (anonymous):

what do you get when you factor out a sin(theta) from that expression in the denominator?

OpenStudy (anonymous):

uhhhh ok how do I factor it out?

OpenStudy (anonymous):

How do you factor out the a from 2a + ba ?

OpenStudy (anonymous):

Or factor out a 2 from (4 + 6a)

OpenStudy (anonymous):

sin(theta)(1/cos(theta)

OpenStudy (anonymous):

close

OpenStudy (anonymous):

try this one: factor completely (10x+2)

OpenStudy (anonymous):

2(5x+1)

OpenStudy (anonymous):

right. Now try \(\frac{a}{b} + a\)

OpenStudy (anonymous):

or we can write it as\[\frac{1}{b}*a + a\]

OpenStudy (anonymous):

if that makes it clearer

OpenStudy (anonymous):

sin(θ)(1+cos(θ))/cos(θ)

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[\frac{a}{b} + a = a(\frac{1}{b} + 1). a = sin(\theta), b = cos(\theta) \]\[\implies \frac{sin(\theta)}{cos(\theta)} + sin(\theta) = sin\theta(\frac{1}{cos\theta} + 1)\]

OpenStudy (anonymous):

oh ok.... i guess i was kinda close, but not really

OpenStudy (anonymous):

so from there you should see something nice to cancel and get a simple answer

OpenStudy (anonymous):

what like the 1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Like the [1 + 1/cos(theta)] you have on top and bottom

OpenStudy (anonymous):

what equation are you looking at?

OpenStudy (anonymous):

\[\frac{1+\frac{1}{cos\theta}}{sin\theta (1 + \frac{1}{cos\theta})}\]

OpenStudy (anonymous):

so all I have left is 1/sin(theta)?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

which equals...csc(theta)

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

wow sorry about making that so difficult thank you so much!

OpenStudy (anonymous):

lol, np